OpenStudy (anonymous):

Item to rent costs $85.00 first 3 hrs, and $19.50 each add. hour, or part of. Budget is $185. How long can item be rented? Just need to equation. thanks

OpenStudy (anonymous):

\[3+\left(\frac{185-85}{19.50}\right)\text{ = }8.12\text{ hours}\]

OpenStudy (anonymous):

cheers :)

OpenStudy (anonymous):

Your welcome.